# Open Source Implementations of Numerical Algorithms for

### v4

@@ Appendix A Error Estimation of Gaussian Numerical Integration
A.1 Gauss-Chebyshev Formulae and Approximation Error
@@
For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}\frac{f(x)}{\sqrt{1-x^{2}}}\operatorname{d}x}$ can be computed by the Gauss-Chebyshev formulae
@@
According to (65) and the $k$-th order derivative formulae
@@
For $f(\cdot,k)\in C^{2n}[-1,1]$ and fixed parameter $k$, the integral ${\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x}$ can be computed with the extended Gauss-Chebyshev formulae
@@
For illustration, if we take $f(x)=\cos x$, then $f^{(2n)}(x)=(-1)^{n}\cos x,f^{(2n)}(kx)=k^{2n}(-1)^{n}\cos kx$. Therefore, the approximation error of integral ${\int^{1}_{-1}\frac{\cos kx}{\sqrt{1-x^{2}}}\operatorname{d}x}=J_{0}(\left|k\right|),\quad k\in\mathbb{R}$ with the Gauss-Chebyshev formulae will be
@@
since $\left|(-1)^{n}\cos(k\xi)\right|\leq 1$. For a fixed $k$, we can set the parameter $n$ as
@@
A.2 Gauss-Legendre Formulae and Approximation Error
@@
For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}f(x)\operatorname{d}x}$ can be computed with the Gauss-Legendre formulae
@@
For $f(\cdot,k)\in C^{2n}[a,b]$ and fixed parameter $k$, the integral ${\int^{b}_{a}f(x,k)\operatorname{d}x}$ can be computed with the extended Gauss-Legendre formulae
@@ Appendix B Orthogonal Polynomials
Table 4: Parameters for the iterative formulae of orthogonal polynomials $Y_{n}(x)$ such that $Y_{0}(x)=c_{0},Y_{1}(x)=c_{1}x+c_{2}$ and $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ for $n\geq 2$.
@@ Appendix D Test Cases for Verification and Validation
$\operatorname{AGM}(a,b)$ $a,b\in\mathbb{R}^{+}$ $\operatorname{AGM}(b,a)$
$\operatorname{AGM}(a,b)$ $a,b\in\mathbb{R}^{+}$ $\operatorname{AGM}\left(\cfrac{a+b}{2},\sqrt{ab}\right)$
@@
  • For computing the root of nonlinear equation $f(x,\alpha)=0$ with Algorithm 3.2, Algorithm 3.2 and Algorithm 3.7, we can construct
    $f(x)=\prod^{n}_{i=1}(x-\gamma_{i})$
    to test the numerical solution to the analytical solution $x_{i}=\gamma_{i}$ for positive $n$ and $\gamma_{1}<\gamma_{2}<\cdots<\gamma_{n}$, say $\gamma_{i}=i$.
@@
    by the Newton-Leibnitz formulae and the rule of L’Hospital in calculus. This result also can be verified with numerical integration.
@@
Code availability statement

### v5

@@ Appendix A Error Estimation of Gaussian Numerical Integration
A.1 Gauss-Chebyshev Formula and Approximation Error
@@
For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}\frac{f(x)}{\sqrt{1-x^{2}}}\operatorname{d}x}$ can be computed by the Gauss-Chebyshev formula
@@
According to (65) and the $k$-th order derivative formula
@@
For $f(\cdot,k)\in C^{2n}[-1,1]$ and fixed parameter $k$, the integral ${\int^{a}_{-a}\frac{f(x,k)}{\sqrt{a^{2}-x^{2}}}\operatorname{d}x}$ can be computed with the extended Gauss-Chebyshev formula
@@
For illustration, if we take $f(x)=\cos x$, then $f^{(2n)}(x)=(-1)^{n}\cos x,f^{(2n)}(kx)=k^{2n}(-1)^{n}\cos kx$. Therefore, the approximation error of integral ${\int^{1}_{-1}\frac{\cos kx}{\sqrt{1-x^{2}}}\operatorname{d}x}=J_{0}(\left|k\right|),\quad k\in\mathbb{R}$ with the Gauss-Chebyshev formula will be
@@
since $\left|(-1)^{n}\cos(k\xi)\right|\leq 1$. For the fixed $k$, we can set the parameter $n$ as
@@
A.2 Gauss-Legendre Formula and Approximation Error
@@
For $f\in C^{2n}[-1,1]$, the integral ${\int^{1}_{-1}f(x)\operatorname{d}x}$ can be computed with the Gauss-Legendre formula
@@
For $f(\cdot,k)\in C^{2n}[a,b]$ and fixed parameter $k$, the integral ${\int^{b}_{a}f(x,k)\operatorname{d}x}$ can be computed with the extended Gauss-Legendre formula
@@ Appendix B Orthogonal Polynomials
Table 4: Parameters for the iterative formula of orthogonal polynomials $Y_{n}(x)$ such that $Y_{0}(x)=c_{0},Y_{1}(x)=c_{1}x+c_{2}$ and $Y_{n}(x)=[A(n)x+B(n)]Y_{n-1}(x)-C(n)Y_{n-2}(x)$ for $n\geq 2$.
@@ Appendix D Test Cases for Verification and Validation
$\operatorname{AGM}(a,b)$ $a,b\in\mathbb{R}^{+}$ $\operatorname{AGM}(b,a)=\operatorname{AGM}\left(\cfrac{a+b}{2},\sqrt{ab}\right)$
@@
  • For computing the root of nonlinear equation $f(x,\alpha)=0$ with Algorithm 3.2, Algorithm 3.2 and Algorithm 3.7, we can construct $f(x)=\prod^{n}_{i=1}(x-\gamma_{i})$ to test the numerical solution to the analytical solution $x_{i}=\gamma_{i}$ for positive $n$ and $\gamma_{1}<\gamma_{2}<\cdots<\gamma_{n}$, say $\gamma_{i}=i$.
@@
    by the Newton-Leibnitz formula and the rule of L’Hospital in calculus. This result also can be verified with numerical integration.
@@
Code & data availability statement